OpenStudy (anonymous):

If a snowball melts so that its surface area decreases at a rate of 1 cm2/min, find the rate at which the diameter decreases when the diameter is 9 cm. (Give your answer correct to 4 decimal places.)

OpenStudy (anonymous):

@Science_ALLY

OpenStudy (anonymous):

you need to equations, one for surface area and one for volume. then you'll need to differentiate with respect to time (each one).

OpenStudy (anonymous):

i think I only need surface area to solve this

OpenStudy (jhannybean):

Okay.

OpenStudy (jhannybean):

A cone is a spherical shape is it not? So whats the surface area of a sphere.

OpenStudy (anonymous):

2pir^(2)

OpenStudy (anonymous):

yeah, just surface area\[A=4 \pi r^2 \Rightarrow \frac{ dA }{ dt }=8 \pi r \frac{ dr }{ dt }\]

OpenStudy (anonymous):

you want \[2\frac{ dr }{ dt }\]

OpenStudy (jhannybean):

And you're looking for the rate at which the diameter decreases. r = \(\frac{D}{2}\)

OpenStudy (anonymous):

\[\frac{ dA }{ dt }=1\frac{ cm^2 }{ \min }\]

OpenStudy (anonymous):

\[\frac{ dD }{dt }=\frac{ dA }{ dt }\cdot \frac{ 1 }{ 2 \pi\cdot 9\,\, cm }\]

OpenStudy (anonymous):

\[A=4\pi r^2 = \pi D^2 \Rightarrow \frac{ dA }{ dt }=2\pi D \frac{ dD }{ dt }\Rightarrow \frac{ dD }{ dt }=\frac{ dA }{ dt }\cdot \frac{ 1 }{ 2\pi D }\]

OpenStudy (anonymous):

I got it!

OpenStudy (anonymous):

maybe sign, try negative of your answer (for decreasing)

OpenStudy (anonymous):

I didn't put parenthesis in the right place in my calculator

OpenStudy (anonymous):

there you go!

OpenStudy (anonymous):

thank you!